7.2 kg object has 95 J of potential energy after being lifted vertically from the floor. How high above the floor is the object? Include a sketch showing the starting position and final position of the object. 

 



1
Expert's answer
2021-06-07T09:29:07-0400


Let's denote the final height as "h". By definition, the potential energy is:


"E = mgh"

where "m = 7.2kg" is the mass of the object, and "g = 9.8N\/kg" is the gravitational acceleration. Expressing "h" and substituting "E = 95J", obtain:


"h = \dfrac{E}{mg}\\
h = \dfrac{95J}{7.2kg\cdot 9.8N/kg} \approx 1.3 m"

Answer. 1.3 m.